A car moves with a speed of 1800km/hr, calculate the distance travelled in half a minute


1
Expert's answer
2021-01-10T18:27:07-0500

Let's first convert km/hr to m/s:


"v=(1800\ \dfrac{km}{hr})\cdot (\dfrac{1000\ m}{1\ km})\cdot(\dfrac{1\ hr}{3600\ s})=500\ \dfrac{m}{s}."

Finally, we can calculate the distance travelled by the car in half a minute:


"d=vt=500\ \dfrac{m}{s}\cdot 30\ s=15000\ m=15\ km."

Answer:

"d=15000\ m=15\ km."